# BioDisco: Multi-agent hypothesis generation with dual-mode evidence, iterative feedback and temporal evaluation

### v1

@@ Results
Analysis
Survey responses were modelled using a single polytomous Rasch model, implemented as a Bayesian cumulative probit mixed effects model (Bürkner 2021) for all four metrics—one for each group of experts:
@@ 5 Discussion & Conclusion
Limitations
Hypotheses were rated as significant and novel but apparently at the expense of relevance and verifiability. One explanation for this could be an inherent trade-off between novelty and verifiability: things that are harder to test are less likely to have already been tested. Additionally, a single LLM judge may not be the best evaluator of experimental feasibility; this is explored further by a dedicated hypothesis testing framework, e.g. Popper (Huang et al. 2025a). Experimental validation in the laboratory is, unfortunately, beyond the scope of this paper.
@@
Future work
A direct comparison of competing hypothesis generation systems, would be a valuable contribution. A significant obstacle may be disentangling the effects of architectural design from more mundane aspects, such as data and knowledge graph availability and the relative strength of different LLMs used in agents or their evaluation. This, like any type of evaluation of LLMs trained on public data, can also be vulnerable to data leakage (see, e.g. Zhou et al. 2023).
@@
Appendices
See Appendix for details on LLM agents and judges, their prompts, gold-standard hypotheses, the human evaluation survey, item response models and computate infrastructure.
@@ Acknowledgments
This work was supported by funding from the German Ministry of Research, Technology and Space (BMFTR) under grant curAIknow 03ZU1202NB, part of project curATime (Cluster for Atherothrombosis and Individualized Medicine; https://curatime.org/).
@@ Appendix A A Detailed Case Study
To provide an insight into how BioDisco works, we present a detailed case study. The case is grounded in a recent study linking G-protein coupled receptors (GPCRs) to increased cardiovascular risk (Kobilka and Lefkowitz 2024), and centres on two entities: the gene GPR153 and the disease vascular injury. We provided “Role of GPR153 in vascular injury and disease” as input to BioDisco. Relevant entities were drawn from the input query and used to access the literature interface to retrieve relevant publications.
@@
“GPR153 activation in vascular smooth muscle cells enhances pro-inflammatory gene expression via the YAP/TAZ pathway, promoting neointima formation following vascular injury”.
Subsequently, the initial hypothesis and relevant keywords were used to query the literature interface for relevant PubMed publications to find evidence for the generated hypothesis. This search identified studies such as (Shao et al. 2025) which conveys how GPR153 is an orphan receptor that facilitates expression of pro-inflammation and pro-proliferation genes in smooth muscle cells by regulating cAMP levels in cells and thereby contributing to inflammation and vascular remodelling.
@@
Following this, the system re-entered the feedback loop. The Critic referencing the newly identified literature for the revised hypothesis, raised the score of Novelty to 5. Elaborating on the mechanism by which GPR153 can influence YAP1 signalling led to an increase in the Significance score to 5 as well. Relevance remained at 5 while Verifiability was still at 4, making the overall score 19. This caused it to access the knowledge graph to retrieve a more focused subgraph containing genes and proteins: GPR153, CEBPB, NRF1, CD7, EGR1, PHYHIP, TTR, PPP2CA, and GSK3B. The system also accessed the literature interface for each major gene and gene combination implicated in the new hypothesis to identify additional relevant publications. This step provided further evidence supporting the possible connections among these genes.
Subsequently, the Reviewer chooses to focus on Verifiability - the only criterion not awarded a top score. The Refiner takes the previous refined hypothesis as input, along with the new subgraph, additional PubMed articles and the Reviewer’s note to generate a further refined hypothesis:
@@
In the next iteration, the Critic assigned scores of 5 for Novelty, Significance, and Relevance. Verifiability remained at 4, as the complexity of the inferred connections could pose challenges for experimental validation. The system regrouped the key genes and conducted an updated literature search, identifying nine unique articles, including some newly retrieved papers in addition to prior results. The Refiner then incorporated additional proteins (NRF1, CD7, and GSK3B) into the hypothesis, yielding the final version:
@@
For this hypothesis, the Critic rated Novelty a 5, highlighting the integration of multiple key regulatory proteins into a comprehensive signaling network. Relevance and Significance also received top scores. Verifiability remained at 4, as the complexity of the interactions may limit definitive experimental validation, despite available methods such as gene editing and pathway analysis. The final hypothesis thus achieved an overall score of 19 out of 20.
Appendix B Feature comparison
CoSci IntS SciA ResA BioDisco
Multi-agent
Tool use
KG
Search
Reviewer
Scoring
Table 2: Features of LLM agentic systems: multiple agents, use of external tools, knowledge graph integration, ability (dynamically) to search academic literature, presence of a ‘reviewer’ agent critically appraising outputs, and refinement of hypotheses based on explicit numerical scoring; CoSci = AI co-scientist (Gottweis et al. 2025); IntS = Intelliscope (Aamer et al. 2025); SciA = SciAgents (Ghafarollahi and Buehler 2024); ResA = ResearchAgent (Baek et al. 2025). ResearchAgent constructs a knowledge graph from an academic graph, but does not access a ‘live’ search API
Appendix C Additional Results
Examples from (Qi et al. 2024)
@@ Appendix C Additional Results
Appendix D Paired Comparison Models
Paired comparison models were fitted using the R package BradleyTerry2 (Turner and Firth 2012), with quasi-variance approximations computed using qvcalc (Firth 2025). The Bradley–Terry model (Bradley and Terry 1952) is given by
$\log\operatorname{odds}(i~\text{beats}~j)=\alpha+\beta_{i}-\beta_{j},$ (4)
where $i$ and $j$ denote ‘players’ or LLM agents, $\beta_{i}$ is the ability score (on the logit-scale) of the $i$th LLM system and $\alpha$ is a ‘home advantage’ parameter, estimating the possible order effect bias, assuming the hypothesis of $i$ is presented to the evaluator first in the pair.
For our model, ties were treated as half wins to each player. An extension of the Bradley–Terry model that explicitly supports ties, the (Davidson 1970) model, was also fitted separately, but yielded very similar results (not presented here).
The Bradley–Terry model parameter estimates $\hat{\alpha}$ are given in Table 6 and suggest a statistically significant order bias for Novelty and Verifiability evaluations (in favour of the item appearing first in the comparisons) but there is not enough evidence to indicate such an effect exists in Relevance or Significance comparisons at the 5% level of significance.
Metric Estimate Std. Err. Statistic $p$-value
Novelty 0.80 0.36 2.23 0.03
Relevance 0.08 0.14 0.54 0.59
Significance 0.07 0.24 0.28 0.78
Verifiability 0.44 0.15 2.95 0.00
Quasi-variances (Firth 2004) aim to minimize the squared loss
@@ Appendix D Paired Comparison Models
where $q_{i}$ is the quasi-variance of player $i$ and $v_{ij}$ is the covariance term (i.e. off-diagonal entries between players $i$ and $j$. The package qvcalc returns relative errors as a measure of the quality of this approximation: the distribution of relative errors is given in Figure 7 as a beeswarm plot (Selby 2020) and appear to be mostly acceptable.
@@
Appendix E Compute Infrastructure
@@
Appendix F Human evaluation
@@
Participants had the right to remain anonymous or to be acknowledged by name in subsequent publications. They were asked to give their institutional affiliation, years of experience, and discretized number of publications in their field of expertise. For each hypothesis, the participant was asked to rate it for novelty, significance, verifiability and relevance (to the given input context), each on a scale from 1–5, where 1 is the lowest and 5 is the highest score. The set of metrics is based on (Qi et al. 2024) and the descriptions were the same as given to LLM evaluators (see Listing 8).
@@
[…]
@@
To quantitatively assess human evaluation scores while accounting for both rater and hypothesis-specific variability, we fit a Bayesian cumulative ordinal mixed-effects model to the ordinal scores assigned by each rater. Let $Y_{ijm}$ denote the ordinal rating (on a scale of 1 to 5) given by rater $i$ to hypothesis $j$ on metric $m$. The model estimates the cumulative probability for each rating category $k\in\{1,2,3,4\}$ as
@@
Using brms (Bürkner 2021) we fit a single model across all metrics to increase statistical power, under the assumption that the rater leniency effects are constant across metrics. This approach pools information across all dimensions for more robust estimation of both rater and hypothesis effects.
Appendix G Agent Roles
Here we describe in detail the role of the individual agents and their respective prompts.
G.1 Planner agent
@@
Listing 1: Planner agent
@@ G.2 Background agent
Listing 2: Background agent
@@ G.3 Explorer agent
Listing 3: Explorer agent
@@ G.4 Scientist agent
Listing 4: Scientist agent
@@ G.5 Critic agent
Listing 5: Critic agent
@@ G.6 Reviewer agent
Listing 6: Reviewer agent
@@ G.7 Refiner agent
Listing 7: Refiner agent
@@ Appendix H LLM Evaluators
Listing 8: Direct evaluator
@@
By contrast, a pairwise evaluator compares two hypotheses at a time, and is asked to say which of them is better according to each of the four criteria. Ties are allowed. A Bradley–Terry model was then fitted to the outputs; see Section D.
Listing 9: Pairwise evaluator
@@
- A if A is clearly superior,
- B if B is clearly superior,
- ”0” if they are equal or difference is unclear.
@@
Novelty: Evaluate the novelty of two scientific hypotheses (A and B) given the user input. For each, assign a novelty score from 0 to 3. 0 means theres no novelty, which indicates that the hypothesis is a paraphrase of the background. 1 means theres slight novelty. 2 means theres moderate novelty. 3 means the hypothesis has strong novelty, which gives new insights beyond the background. Score two hypotheses and compare which one is more novel(”A”, B”, or ”0” if equal or difference is unclear)
Relevance: Evaluate the relevance of two scientific hypotheses (A and B) given the user input. For each, assign a relevance score from 0 to 3. 0 means theres no relevance. 1 means theres slight relevance. 2 means theres moderate relevance. 3 means the hypothesis is strongly related to the background. Score both hypotheses and compare which one is more relevant (”A”, B”, or ”0” if equal or difference is unclear)
Significance: Evaluate the significance of two scientific hypotheses (H_A and H_B) given the user input. For each, assign a significance score from 0 to 3. 0 means theres no significance, which indicates that the hypothesis is just common knowledge. 1 means theres slight significance. 2 means theres moderate significance. 3 means the hypothesis has strong significance, providing significant insights beyond the background. Score both hypotheses and compare which one is more significant (”A”, B”, or ”0” if equal or difference is unclear)
Verifiability: Evaluate the verifiability of two scientific hypotheses (H_A and H_B) given the user input. For each, assign a verifiability score from 0 to 3. 0 means theres no verifiability, which indicates that the hypothesis is not possible to be verified in future work. 1 means theres slight verifiability. 2 means theres moderate verifiability. 3 means the hypothesis has strong verifiability, which means it is very likely to be verified in future work. Score both hypotheses and compare which one is more verifiable (”A”, B”, or ”0” if equal or difference is unclear)
@@
  • N. Aamer, M. N. Asim, S. Munir, and A. Dengel (2025) Automating AI discovery for biomedicine through knowledge graphs and LLM agents. bioRxiv. External Links: Document, Link, https://www.biorxiv.org/content/early/2025/05/13/2025.05.08.652829.full.pdf Cited by: Table 2.
  • J. Baek, S. K. Jauhar, S. Cucerzan, and S. J. Hwang (2025) ResearchAgent: iterative research idea generation over scientific literature with large language models. In Proceedings of the 2025 Conference of the Nations of the Americas Chapter of the Association for Computational Linguistics: Human Language Technologies (Volume 1: Long Papers), L. Chiruzzo, A. Ritter, and L. Wang (Eds.), Albuquerque, New Mexico, pp. 6709–6738. External Links: Link, Document, ISBN 979-8-89176-189-6 Cited by: Table 2.
  • R. A. Bradley and M. E. Terry (1952) Rank analysis of incomplete block designs: the method of paired comparisons. Biometrika 39 (3-4), pp. 324–345. External Links: ISSN 0006-3444, Document, Link, https://academic.oup.com/biomet/article-pdf/39/3-4/324/930466/39-3-4-324.pdf Cited by: Appendix D.
  • P. Bürkner (2021) Bayesian item response modeling in r with brms and stan. Journal of statistical software 100, pp. 1–54. Cited by: Appendix F.
  • R. R. Davidson (1970) On extending the bradley-terry model to accommodate ties in paired comparison experiments. Journal of the American Statistical Association 65 (329), pp. 317–328. External Links: ISSN 1537-274X, Link, Document Cited by: Appendix D.
  • D. Firth (2004) Quasi-variances. Biometrika 91 (1), pp. 65–80. External Links: ISSN 1464-3510, Link, Document Cited by: Appendix D.
  • D. Firth (2025) qvcalc: quasi variances for factor effects in statistical models. Note: R package version 1.0.4 External Links: Link, Document Cited by: Appendix D.
  • A. Ghafarollahi and M. J. Buehler (2024) SciAgents: automating scientific discovery through bioinspired multi-agent intelligent graph reasoning. Advanced Materials, pp. 2413523. External Links: Document, https://advanced.onlinelibrary.wiley.com/doi/pdf/10.1002/adma.202413523 Cited by: Table 2.
  • J. Gottweis, W. Weng, A. Daryin, T. Tu, A. Palepu, P. Sirkovic, A. Myaskovsky, F. Weissenberger, K. Rong, R. Tanno, K. Saab, D. Popovici, J. Blum, F. Zhang, K. Chou, A. Hassidim, B. Gokturk, A. Vahdat, P. Kohli, Y. Matias, A. Carroll, K. Kulkarni, N. Tomasev, Y. Guan, V. Dhillon, E. D. Vaishnav, B. Lee, T. R. D. Costa, J. R. Penadés, G. Peltz, Y. Xu, A. Pawlosky, A. Karthikesalingam, and V. Natarajan (2025) Towards an AI co-scientist. External Links: 2502.18864, Document, Link Cited by: Table 2.
@@
  • B. Qi, K. Zhang, K. Tian, H. Li, Z. Chen, S. Zeng, E. Hua, H. Jinfang, and B. Zhou (2024) Large language models as biomedical hypothesis generators: a comprehensive evaluation. In First Conference on Language Modeling, External Links: Link Cited by: Appendix C, Appendix F.
  • D. A. Selby (2020) Statistical modelling of citation networks, research influence and journal prestige. Ph.D. Thesis, University of Warwick. External Links: Link Cited by: Appendix D.
@@
  • H. Turner and D. Firth (2012) Bradley–Terry models in R: the BradleyTerry2 package. Journal of Statistical Software 48 (9), pp. 1–21. External Links: Document Cited by: Appendix D.

### v2

@@ 3.3.1 Methodology
3.3.2 Analysis
Survey responses were modelled using a single polytomous Rasch model, implemented as a Bayesian cumulative probit mixed effects model (Bürkner, 2021) for all four metrics—one for each group of experts:
@@ 3.3.2 Analysis
$\eta_{ijm}=\underbrace{\beta_{m}}_{\text{metric effect}}+\underbrace{u_{i}}_{\text{rater effect}}+\underbrace{v_{jm},}_{\text{hypothesis-on-metric effect}}$ (3)
@@ 5 Discussion & Conclusion
5.1 Limitations
Hypotheses were rated as significant and novel but apparently at the expense of relevance and verifiability. One explanation for this could be an inherent trade-off between novelty and verifiability: things that are harder to test are less likely to have already been tested. Additionally, a single LLM judge may not be the best evaluator of experimental feasibility; this is explored further by a dedicated hypothesis testing framework, e.g. Popper (Huang et al., 2025a). Experimental validation in the laboratory is, unfortunately, beyond the scope of this paper.
@@ 5.1 Limitations
In our TruthHypo evaluation, we excluded the ‘no relation’ class to focus on the system’s ability to distinguish positive from negative associations. However, in a real-world setting, the ability to abstain may prevent hallucinations. Currently, BioDisco incorporates no such low-confidence filter.
5.2 Future work
While our evaluation demonstrates BioDisco’s superior novelty and significance to generalist agents like Biomni (see Figure 4), the comparison does not necessarily disentangle effects of architectural design from more mundane aspects, such as data and knowledge graph availability and the relative strength of different LLMs used in agents or their evaluation. This, like any type of evaluation of LLMs trained on public data, can also be vulnerable to data leakage (see, e.g. Zhou et al., 2023). Furthermore, a purely in silico comparison may only test semantic plausibility; a real-world wet lab evaluation would definitively validate the utility of the system.
@@ Acknowledgments
This work was supported by funding from the German Ministry of Research, Technology and Space (BMFTR) under grant curAIknow 03ZU1202NB, part of project curATime (Cluster for Atherothrombosis and Individualized Medicine; https://curatime.org).
@@ Appendix A A Detailed Case Study
To provide a robust and unbiased assessment of BioDisco’s generated hypotheses, we conducted a structured case study combining automated outputs with expert review. The case is grounded in a recent study linking G-protein coupled receptors (GPCRs) to increased cardiovascular risk (Kobilka and Lefkowitz, 2024), and centres on two entities: the gene GPR153 and the disease vascular injury. We provided “Role of GPR153 in vascular injury and disease” as input to BioDisco. Relevant entities were drawn from the input query and used to access the literature interface to retrieve relevant publications.
@@
“GPR153 activation in vascular smooth muscle cells enhances pro-inflammatory gene expression via the YAP/TAZ pathway, promoting neointima formation following vascular injury”
Subsequently, the initial hypothesis and relevant keywords were used to query the literature interface for relevant PubMed publications to find evidence for the generated hypothesis. This search identified studies such as Shao et al. (2025) which conveys how GPR153 is an orphan receptor that facilitates expression of pro-inflammation and pro-proliferation genes in smooth muscle cells by regulating cAMP levels in cells and thereby contributing to inflammation and vascular remodelling.
@@
Following this, the system re-entered the feedback loop. Incorporating newly identified literature, the Critic raised the Novelty and Significance scores to 5, while Relevance remained at 5 and Verifiability at 4, yielding an overall score of 19. Normally, iteration would stop here, as further improvement is constrained by the inherent complexity of the biology rather than lack of evidence. For demonstration, however, we continued the process to illustrate how the system explores more intricate mechanistic hypotheses. Additional literature searches for each key gene and their combinations further substantiated the potential connections identified in the evolving hypothesis. Subsequently, the Reviewer focused on Verifiability—the only criterion not awarded a top score. Incorporating new subgraph information and additional literature, the Refiner further elaborated the hypothesis to include a broader CEBPB-mediated network involving YAP1, EGR1, and GSK3B:
@@
In the next round, the Critic again assigned top scores for Novelty, Significance, and Relevance, while Verifiability remained at 4, reflecting the growing experimental challenges posed by increased mechanistic complexity. To further demonstrate system capabilities, additional genes (NRF1, CD7, and GSK3B) were integrated, supported by targeted literature searches:
@@
For this, Critic gave Novelty a 5 because it "combines several key regulatory proteins" to provide a comprehensive signalling network. Both Relevance and Significance were also scored at 5. Verifiability was still at 4 because Critic noted that although techniques like gene editing, pathway analysis and in vivo models of vascular injury could be used to test the hypothesis, the "complexity of the interactions may pose challenges in definitively confirming all proposed relationships". This resulted in an overall score of 19/20.
Appendix B Additional Results
Examples from Qi et al. (2024)
@@ Appendix B Additional Results
Appendix C Paired Comparison Models
Paired comparison models were fitted using the R package BradleyTerry2 (Turner and Firth, 2012), with quasi-variance approximations computed using qvcalc (Firth, 2025). The Bradley–Terry model (Bradley and Terry, 1952) is given by
$\log\operatorname{odds}(i~\text{beats}~j)=\beta_{i}-\beta_{j}+\alpha z_{ij},$ (4)
where $i$ and $j$ denote ‘players’ or LLM configurations, $\beta_{i}$ is the ability score of the $i$th LLM system and $\alpha$ is a ‘home advantage’ parameter, estimating the possible order effect bias, with $z_{ij}=1$ if the hypothesis of $i$ is presented to the evaluator first in the pair, and $z_{ij}=-1$ otherwise.
For our model, ties were awarded as half wins to each player. An extension of the Bradley–Terry model that explicitly supports ties, the Davidson (1970) model, was also fitted separately, but yielded very similar results (not presented here).
Quasi-variances (Firth, 2004) aim to minimize the squared loss
@@ Appendix C Paired Comparison Models
where $q_{i}$ is the quasi-variance of player $i$ and $v_{ij}$ is the covariance term (i.e. off-diagonal entries between players $i$ and $j$. The package qvcalc returns relative errors as a measure of the quality of this approximation: the distribution of relative errors is given in Figure 7 as a beeswarm plot (Selby, 2020) and appear to be mostly acceptable.
@@
Appendix D Compute Infrastructure
@@
Appendix E Human evaluation
@@
Participants had the right to remain anonymous or to be acknowledged by name in subsequent publications. They were asked to give their institutional affiliation, years of experience, and discretized number of publications in their field of expertise. For each hypothesis, the participant was asked to rate it for novelty, significance, verifiability and relevance (to the given input context), each on a scale from 1–5, where 1 is the lowest and 5 is the highest score. The set of metrics is based on Qi et al. (2024) and the descriptions were the same as given to LLM evaluators (see Prompt 8).
@@
PTLOs and immune responses in atherosclerosis Activation of PTGDS within tertiary lymphoid organ-like structures enhances B cell-mediated antibody production, promoting plaque instability in atherosclerosis. Activation of PTGDS within tertiary lymphoid organ-like structures not only enhances B cell-mediated antibody production but also establishes a feedback loop involving TNF-$\alpha$ and AKT1 that regulates macrophage efferocytosis, thereby orchestrating inflammatory responses that escalate plaque instability in atherosclerosis.
@@
A summary of this model appears in § 3.3 (“What do human domain experts think?”) of the main text, and we provide the full details here. To quantitatively assess human evaluation scores while accounting for both rater and hypothesis-specific variability, we fit a Bayesian cumulative ordinal mixed-effects model to the ordinal scores assigned by each rater. Let $Y_{ijm}$ denote the ordinal rating (on a scale of 1 to 5) given by rater $i$ to hypothesis $j$ on metric $m$. The model estimates the cumulative probability for each rating category $k\in\{1,2,3,4\}$ as
@@
Using brms (Bürkner, 2021) we fit a single model across all metrics to increase statistical power, under the assumption that the rater leniency effects are constant across metrics. This approach pools information across all dimensions for more robust estimation of both rater and hypothesis effects.
Appendix F Agent Roles
Here we describe in detail the role of the individual agents, their respective prompts and the technical implementation of interactions with the knowledge graph and PubMed.
F.1 Implementation Details
Adaptive adjustment of KG query parameters
First, the upstream agent (e.g. ReviewerAgent) automatically identifies suboptimal evaluation criteria based on task objectives and intermediate feedback, and employs LLM-based reasoning to generate updated traversal depth (DEPTH_OVERRIDE) and relation type lists (RELS_OVERRIDE). Second, the downstream KG interface merges default relation types and maximum traversal depth for the specified domain of the hypothesis (e.g. the molecular domain includes protein_protein, molfunc_protein and bioprocess_protein relations by default). If override parameters are provided, they fully replace the defaults, directly controlling both single-hop and multi-hop KG retrieval. If the result exceeds a predefined threshold, the system automatically reverts to the domain default settings and applies top-$k$ pruning. This mechanism ensures that relation types and traversal depth are adaptively optimized across domains in response to task requirements and real-time feedback.
Standardized representation of KG subgraphs
The KG interface returns subgraphs as structured JSON objects containing nodes, direct edges, and multi-hop paths. Each node entry includes a unique identifier, name, type, and standardized attributes, while each edge specifies the source, target, relation type, weight and path information. For downstream use, the system summarizes and converts these subgraphs into standardized plaintext: each node is represented as ‘name(type)’, each direct edge as ‘source(type) → relation type(weight) → target(type)’ and each multi-hop path as a sequential concatenation of nodes and relations. This process produces concise, human-readable text formats suitable for subsequent LLM-based analysis or other downstream tasks.
LLM-based PubMed querying strategy
We employ LLM-powered agents to analyze user-provided keywords, hypotheses, and associated feedback. Through carefully designed system prompts, the LLM automatically groups semantically similar keywords into synonym sets (using OR logic within each group), and determines the optimal Boolean logic (AND or OR) for combining these groups based on context and query intent. The LLM returns a structured JSON output specifying grouped terms, Boolean logic and multiple candidate query formulations with guidance notes.
The system then parses this structured output and generates PubMed API-compatible query strings via templated rules: each group of terms is tagged for MeSH or TIAB fields, inter-group logic is set as recommended by the LLM, and query constraints such as date range and publication type are automatically added. For each generated query, the system executes PubMed searches in sequence; if insufficient results are obtained, subgroup queries are recursively constructed. All retrieved articles are deduplicated, merged and ranked by publication date. This workflow enables robust, context-adaptive PubMed retrieval by tightly integrating LLM-driven planning with programmatic query construction and post-processing.
F.2 Planner agent
@@
Prompt 1: Planner agent
@@ F.3 Background agent
Prompt 2: Background agent
@@ F.4 Explorer agent
Prompt 3: Explorer agent
@@ F.5 Scientist agent
Prompt 4: Scientist agent
@@ F.6 Critic agent
Prompt 5: Critic agent
@@ F.7 Reviewer agent
Prompt 6: Reviewer agent
@@ F.8 Refiner agent
Prompt 7: Refiner agent
@@ Appendix G LLM Evaluators
Prompt 8: Direct evaluator
@@
By contrast, a pairwise evaluator compares two hypotheses at a time, and is asked to say which of them is better according to each of the four criteria. Ties are allowed. A Bradley–Terry model was then fitted to the outputs; see Section C.
Prompt 9: Pairwise evaluator
@@
- "A" if A is clearly superior,
- "B" if B is clearly superior,
- "0" if they are equal or difference is unclear.
@@
Novelty: Evaluate the novelty of two scientific hypotheses (A and B) given the user input. For each, assign a novelty score from 0 to 3. 0 means theres no novelty, which indicates that the hypothesis is a paraphrase of the background. 1 means theres slight novelty. 2 means theres moderate novelty. 3 means the hypothesis has strong novelty, which gives new insights beyond the background. Score two hypotheses and compare which one is more novel("A", "B", or "0" if equal or difference is unclear)
Relevance: Evaluate the relevance of two scientific hypotheses (A and B) given the user input. For each, assign a relevance score from 0 to 3. 0 means theres no relevance. 1 means theres slight relevance. 2 means theres moderate relevance. 3 means the hypothesis is strongly related to the background. Score both hypotheses and compare which one is more relevant ("A", "B", or "0" if equal or difference is unclear)
Significance: Evaluate the significance of two scientific hypotheses (H_A and H_B) given the user input. For each, assign a significance score from 0 to 3. 0 means theres no significance, which indicates that the hypothesis is just common knowledge. 1 means theres slight significance. 2 means theres moderate significance. 3 means the hypothesis has strong significance, providing significant insights beyond the background. Score both hypotheses and compare which one is more significant ("A", "B", or "0" if equal or difference is unclear)
Verifiability: Evaluate the verifiability of two scientific hypotheses (H_A and H_B) given the user input. For each, assign a verifiability score from 0 to 3. 0 means theres no verifiability, which indicates that the hypothesis is not possible to be verified in future work. 1 means theres slight verifiability. 2 means theres moderate verifiability. 3 means the hypothesis has strong verifiability, which means it is very likely to be verified in future work. Score both hypotheses and compare which one is more verifiable ("A", "B", or "0" if equal or difference is unclear)
@@
  • R. A. Bradley and M. E. Terry (1952) Rank analysis of incomplete block designs: the method of paired comparisons. Biometrika 39 (3-4), pp. 324–345. External Links: ISSN 0006-3444, Document, Link, https://academic.oup.com/biomet/article-pdf/39/3-4/324/930466/39-3-4-324.pdf Cited by: Appendix C.
  • P. Bürkner (2021) Bayesian item response modeling in r with brms and stan. Journal of statistical software 100, pp. 1–54. Cited by: Appendix E.
  • R. R. Davidson (1970) On extending the bradley-terry model to accommodate ties in paired comparison experiments. Journal of the American Statistical Association 65 (329), pp. 317–328. External Links: ISSN 1537-274X, Link, Document Cited by: Appendix C.
  • D. Firth (2004) Quasi-variances. Biometrika 91 (1), pp. 65–80. External Links: ISSN 1464-3510, Link, Document Cited by: Appendix C.
  • D. Firth (2025) qvcalc: quasi variances for factor effects in statistical models. Note: R package version 1.0.4 External Links: Link, Document Cited by: Appendix C.
@@
  • B. Qi, K. Zhang, K. Tian, H. Li, Z. Chen, S. Zeng, E. Hua, H. Jinfang, and B. Zhou (2024) Large language models as biomedical hypothesis generators: a comprehensive evaluation. In First Conference on Language Modeling, External Links: Link Cited by: Appendix B, Appendix E.
  • D. A. Selby (2020) Statistical modelling of citation networks, research influence and journal prestige. Ph.D. Thesis, University of Warwick. External Links: Link Cited by: Appendix C.
@@
  • H. Turner and D. Firth (2012) Bradley–Terry models in R: the BradleyTerry2 package. Journal of Statistical Software 48 (9), pp. 1–21. External Links: Document Cited by: Appendix C.